# Deviation Tests for a High-dimensional Mean

### 15 Mar 2026 v1

@@ 2 One-sample case
$\theta_{t}=\left\{\begin{array}[]{cc}1,&\text{if }\ \mathcal{M}_{t-1,T_{2}}(\vec{\theta}_{t-1})\leq 0,\\[6.0pt] -1,&\text{if }\ \mathcal{M}_{t-1,T_{2}}(\vec{\theta}_{t-1})>0.\end{array}\right$.
@@
where $\mathbb{E}[X_{t}]=\|\boldsymbol{\mu}\|_{2}^{2}-d_{0}^{2}$ by (9). Intuitively, under $H_{0}$, $\mathbb{E}[X_{t}]\geq 0$, and the parameter $\theta_{t}$ makes the innovation $\tilde{h}(X_{t})$ to have in average an opposite sign to the current value $\mathcal{M}_{t-1,T_{2}}(\vec{\theta}_{t-1})\leq 0$ and vice versa, which defines a negative feedback process. By contrast, $\{\mathcal{M}_{t,T_{2}}(\vec{\theta}_{t})\}$ acts as a positive feedback under $H_{1}$, where $\mathbb{E}[X_{t}]<0$. Consequently, as $t\to\infty$, the negative feedback process will restrict our statistic $\mathcal{M}_{t,T_{2}}(\vec{\theta}_{t})$ asymptotically concentrating around 0 under $H_{0}$, while the positive feedback process will push $\mathcal{M}_{t,T_{2}}(\vec{\theta}_{t})$ asymptotically diverging to infinity under $H_{1}$.
Actually, $\big\{\mathcal{M}_{t,T_{2}}(\vec{\theta}_{t})\big\}$ follows a TAB process and $\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})$ asymptotically converges to the bandit distribution $\mathcal{B}(\kappa)$, whose density function has the form of

### 19 Mar 2026 v2

@@ 2.1 Test statistic
$\theta_{t}=\left\{\begin{array}[]{cc}1,&\text{if }\ \mathcal{M}_{t-1,T_{2}}(\vec{\theta}_{t-1})\leq 0,\\[6.0pt] -1,&\text{if }\ \mathcal{M}_{t-1,T_{2}}(\vec{\theta}_{t-1})>0.\end{array}\right$. (13)
@@
where $\mathbb{E}[X_{t}]=\|\boldsymbol{\mu}\|_{2}^{2}-d_{0}^{2}$ by (9). Intuitively, under $H_{0}$, $\mathbb{E}[X_{t}]\geq 0$, and the parameter $\theta_{t}$ makes the innovation $\tilde{h}(X_{t})$ to have in average an opposite sign to the current value $\mathcal{M}_{t-1,T_{2}}(\vec{\theta}_{t-1})\leq 0$ and vice versa, which defines a negative feedback process. By contrast, $\{\mathcal{M}_{t,T_{2}}(\vec{\theta}_{t})\}$ acts as a positive feedback under $H_{1}$, where $\mathbb{E}[X_{t}]<0$. Hence, as $t\to\infty$, the negative feedback process will restrict our statistic $\mathcal{M}_{t,T_{2}}(\vec{\theta}_{t})$ asymptotically concentrating around 0 under $H_{0}$, while the positive feedback process will push $\mathcal{M}_{t,T_{2}}(\vec{\theta}_{t})$ asymptotically diverging to infinity under $H_{1}$.
2.2 Rationale of the test statistic
There is important difference between our test statistic $\mathcal{M}_{t,T_{2}}(\vec{\theta}_{t})$ in (10) and the one for the univariate deviation test in [6], the former having little in common with the latter. Precisely, given the independent univariate samples $\{x_{t}\in\mathbb{R}:t=1,\cdots,T\}$ with population mean of $\mu>0$, [6] consider the following univariate deviation test:
$H_{0}^{\texttt{u}}:\mu-d_{0}>0\quad\text{versus}\quad H_{1}^{\texttt{u}}:\mu-d_{0}\leq 0.$ (15)
[6] constructed their statistic $\hat{b}_{T,t}^{\texttt{u}}$ as in (5), where $\hat{b}_{T,1}^{\texttt{u}}=x_{1}-d_{0}$ and for $t=2,\cdots,T$
$\hat{b}_{T,t}^{\texttt{u}}=\hat{b}_{T,t-1}^{\texttt{u}}+h(x_{t}),\quad h(x_{t})=\eta_{t}(x_{t}-d_{0}),\quad\eta_{t}=\left\{\begin{array}[]{cc}1,&\text{if }\ \hat{b}_{T,t-1}^{\texttt{u}}\leq 0,\\[6.0pt] -1,&\text{if }\ \hat{b}_{T,t-1}^{\texttt{u}}>0.\end{array}\right$. (18)
By our previous discussion, $\hat{b}_{T,t}^{\texttt{u}}$ acts as a negative/positive feedback process under $H_{0}/H_{1}$. Notably, to derive such properties, the innovation function $h(\cdot)$ is a product of the control $\eta_{t}$ and the target variable $\mathcal{T}_{t}:=x_{t}-d_{0}$, where $\eta_{t}$ controls the sign of $\mathcal{T}_{t}$ based on current value $\hat{b}_{T,t-1}^{\texttt{u}}$ and $\mathbb{E}[\mathcal{T}_{t}]=\mu-d_{0}$ is the term to test in (15). Particularly, the independence between $\mathcal{T}_{t}$ and $\hat{b}_{T,t-1}^{\texttt{u}}$ is important in technical proofs in [6].
[…]
$\hat{b}_{T,t}^{\texttt{h}}=\hat{b}_{T,t-1}^{\texttt{h}}+\theta_{t}\tilde{\mathcal{T}}_{t},\quad\tilde{\mathcal{T}}_{t}=\boldsymbol{x}_{t}^{\prime}\boldsymbol{x}_{t}-\operatorname{Tr}(\boldsymbol{\Sigma})-d_{0}^{2},$
Obviously, $\mathbb{E}[\tilde{\mathcal{T}}_{t}]=\|\boldsymbol{\mu}\|_{2}^{2}-d_{0}^{2}$ and $\tilde{\mathcal{T}}_{t}$ is independent of $\hat{b}_{T,t-1}^{\texttt{h}}$, so $\hat{b}_{T,t}^{\texttt{h}}$ follows a negative/positive feedback process as expected. However, even though $\operatorname{Tr}(\boldsymbol{\Sigma})$ is known, $\tilde{\mathcal{T}}_{t}$ faces a main drawback that $\operatorname{Var}(\tilde{\mathcal{T}}_{t})\asymp O(n)$ compared with $\mathbb{E}[\tilde{\mathcal{T}}_{t}]\leq O(1)$. Therefore, the huge variance of $\tilde{\mathcal{T}}_{t}$ will complete mask any significance of the testing term $\mathbb{E}[\tilde{\mathcal{T}}_{t}]=\|\boldsymbol{\mu}\|_{2}^{2}-d_{0}^{2}$ and makes the resulting test powerless.
Therefore, a suitable target variable $\tilde{\mathcal{T}}_{t}$ should simultaneously satisfy $\mathbb{E}[\tilde{\mathcal{T}}_{t}]=\|\boldsymbol{\mu}\|_{2}^{2}-d_{0}^{2}$ and $\operatorname{Var}(\tilde{\mathcal{T}}_{t})\asymp O(1)$. In high-dimensional statistical literature, a widely-used estimator of $\|\boldsymbol{\mu}\|_{2}^{2}$ is $\frac{1}{T(T-1)}\sum_{s\neq t}^{T}\boldsymbol{x}_{s}^{\prime}\boldsymbol{x}_{t}$, see [4]. Inspired by this estimator, we consider $X_{t}=\frac{1}{T_{1}}\sum_{s=1}^{T_{1}}(\boldsymbol{x}_{s}^{\prime}\boldsymbol{x}_{t+T_{1}}-d_{0}^{2})$ in (9) as the target variable, we can show that $\mathbb{E}[X_{t}]=\|\boldsymbol{\mu}\|_{2}^{2}-d_{0}^{2}$ and $\operatorname{Var}(X_{t})\asymp O(1)$.
[…]
2.3 Asymptotic distribution
Asymptotically, $\big\{\mathcal{M}_{t,T_{2}}(\vec{\theta}_{t})\big\}$ follows a TAB process and $\mathcal{M}_{T_{2},T_{2}}(\vec{\theta}_{T_{2}})$ weakly converges to the bandit distribution $\mathcal{B}(\kappa)$, whose density function has the form of